# The Illusion of State in State-Space Models

### v1

@@ 4.3 Discussion
Theorems 4.1 and 4.2 establish that SSMs, like transformers, can only express solutions to problems in the class $\mathsf{TC}^{0}$. This means that SSMs cannot solve $\mathsf{NC}^{1}$-hard problems like evaluating boolean formulas or graph connectivity. In particular, it shows that they are limited as far as their state tracking capabilities as they are unable to compose permutations (solve the $S_{5}$ word problem):
@@ 5 Extending the Expressive Power of SSMs
We have shown that S4 and S6, despite their seemingly “stateful” design, cannot express problems outside $\mathsf{TC}^{0}$, which includes state tracking like $S_{5}$. We show how SSMs can be extended to close the gap in expressive power with RNNs, allowing them to express $S_{5}$. Two simple extensions can bring about this increase in expressive power: adding a nonlinearity to make the SSM more like an RNN or allowing the $A$ matrix to be input-dependent to make the SSM more like a weighted finite automaton (WFA; Mohri, 2009).
@@ 5.1 Via Nonlinearities
Concretely, an RNN-SSM layer with a step activation function can be defined as follows:
$h_{i}=\mathrm{sgn}\left(Ah_{i-1}+Bx_{i}\right).$ (5)
After this change, the SSM no longer has a straightforward convolutional form. However, its recurrent form is effectively an RNN, and can therefore solve $S_{5}$:
@@
For any regular language $L$ (including the word problem for $S_{5}$), there exists a one-layer log-precision RNN-SSM that recognizes $L$.
@@
The standard constructions for simulating automata with RNNs (cf. Minsky, 1954; Merrill, 2019) apply. ∎
Note that adding nonlinearities to the output of an SSM layer (as in Mamba) is not the same thing as an RNN-SSM. Rather, an RNN-SSM has nonlinearities applied after each recurrent update to the state.
@@ 5.2 Via Input-Dependent Transition Matrices
Another completely different way to get greater expressive power is to let $A$ matrix to be input-dependent. To illustate this, we define and analyze the WFA-SSM layer. Let $A(x_{i})=I+s_{A}(x_{i})$. The recurrent form becomes:
$h_{i}=A(x_{i})h_{i-1}+Bx_{i}.$ (6)
This means the convolutional form computes an iterated product of a sequence of matrices rather than powering a matrix as for S4 and S6 (cf. Section 2.1):
$h_{i}=\sum_{j=1}^{i}\left(\prod_{k=j+1}^{i}A(x_{k})\right)Bx_{j}.$ (7)
Unlike matrix powers, iterated matrix products cannot, in general, be computed in $\mathsf{TC}^{0}$ (Mereghetti & Palano, 2000). This means that the argument from Theorem 4.1 will not go through for WFA-SSMs. In fact, we can show that WFA-SSMs gains expressive power beyond $\mathsf{TC}^{0}$:
@@
For any regular language $L$ over vocabulary $\Sigma$ (including the word problem for $S_{5}$), there exists a one-layer log-precision WFA-SSM that recognizes \$$\$L$, where \$$\$\not\in\Sigma$ is a special beginning-of-string symbol.
@@
It suffices to show that an WFA-SSM can simulate a deterministic finite automaton (DFA). We do this via a transition monoid construction. For any $w\in\Sigma^{*}$, let $\delta_{w}:Q\to Q$ be the function mapping a state to its eventual destination state after $w$ is read from that state. For any DFA, this set of functions forms a finite monoid (the transition monoid) under composition, following from the Myhill-Nerode theorem (Hopcroft et al., 2001). Further, each monoid element $\delta_{w}$ can be represented as a boolean transition matrix, making matrix multiplication isomorphic to monoid composition.
Computing the transition monoid of a DFA allows recognizing valid words: compute the monoid element for a word by multiplying the elements for its tokens and then check whether the initial state maps to an accepting state. In fact, a standard way to solve monoid word problems (e.g., for $S_{5}$) with a DFA is simply to construct a DFA whose transition monoid is the monoid of interest.
Fix a DFA and its transition monoid $\delta$. To complete the proof, we show that there exists an SSM that, for all $w\in\Sigma^{*}$, computes $\delta_{w}$ given input $x=\$w$. We view indices in $h_{i}$ as states and define $B\$$ as $1$ at each accepting state, and 00 elsewhere. For all other $\sigma$, we let $B\sigma=\vec{0}$. This reduces the convolutional form of $h_{i}$ to have a single term:
$h_{i}=\prod_{k=2}^{i}A(x_{k}).$
Now, let $A(\sigma)=\delta_{\sigma}$ for $\sigma\in\Sigma$. It follows that
$h_{i}=\prod_{k=2}^{i}A(x_{k})=\prod_{k=2}^{i}\delta_{x_{k}}.$
Since $x=\$w$, we conclude that $h_{\left\lvert x\right\rvert}$ is $\delta_{w}$. ∎
@@ 5.3 Discussion
Parallelism.
To be used effectively in an LLM, a model architecture must be parallelizable on practical hardware. Architectures in $\mathsf{TC}^{0}$ are parallelizable by design (Merrill & Sabharwal, 2023a), but architectures in $\mathsf{NC}^{1}$ may still be parallelizable to log depth even if they cannot be parallelized to constant depth. For the WFA-SSM, the bottleneck would be computing iterated matrix product with a log-depth computation graph. Similarly, for the RNN-SSM, the bottleneck would be computing the state updates with a log-depth binary tree rather than left to right. If this could be accomplished on practical hardware, these architectures could be parallelizable enough to scale on modern hardware.
Learning Dynamics.
Another potential concern for these SSM variants compared to the original SSM is whether the learning dynamics are as good. In particular, for the WFA-SSM, an iterated product of matrices may lead to vanishing gradient issues. However, this is already potentially an issue for the S6 architecture, where the selective gating involves computing an iterated product of scalars.
@@ 6 Can SSMs Learn Permutations in Practice?
Figure 3: Minimum number of layers required to attain $>90\%$ validation accuracy on group multiplication problems by sequence length and group. RNN and IDS4 models of constant depth can solve arbitrarily long sequences, while transformer, S4, and Mamba models require depths monotonically increasing in sequence length.
@@
Having established theoretical limitations of SSMs for state tracking, we now empirically test how well SSMs can learn such tasks, focusing on the word problem for $A_{5}$. Since this problem is $\mathsf{NC}^{1}$-complete and both transformers and SSMs can only express functions in $\mathsf{TC}^{0}$, to solve instances of this problem these models should require a dynamic depth that grows with the input length.
Setup.
@@
We model word problems (see Section 3.1) as a token-tagging task. Models are given as input a sequence of elements $g_{0}g_{1}g_{2}\cdots g_{n}$ drawn from one of $A_{5}$, $A_{4}\times\mathbb{Z}_{5}$, or $\mathbb{Z}_{60}$, and for each step $t_{i}$ must predict the result of multiplying the first $i$ elements of the sequence together. Modeling the problem as a tagging task rather than as a sequence classification task provides the models with more supervision during training, making it as easy as possible to learn the correct function. We tokenize inputs such that each element gets a unique token. We train models on sequences of length $n$ for successively larger values of $n$ and report full-sequence accuracy on a validation set. [^8] We always include the full set of $3600$ pairwise sequences of length $2$ in the training data, along with the training split of length-$n$ sequences. To validate the predictions of SSMs and transformers being depth-bounded for expressing the word problem for $A_{5}$, we sweep over the number of layers each S4, Mamba and transformer model has, and compare these results to a single-layer simple RNN and our proposed single-layer Input-Dependent S4 (IDS4) architecture.
Results.
@@
Figure 3 shows that single-layer RNN and IDS4 models learn the word problem for arbitrarily long sequences from all three groups. In contrast, transformer, S4, and Mamba models require depth monotonically increasing in sequence length to attain good accuracy on a validation set on non-commutative groups. We draw three main conclusions from this:
  1. 1.
@@
    Mamba and S4 show the same qualitative limitations as transformers on the inherently sequential task $A_{5}$: longer $A_{5}$ sequences require deeper models. This is consistent with both transformers, S4, and Mamba being in $\mathsf{TC}^{0}$. On the other hand, RNNs (which resemble automata; Merrill, 2019) and IDS4 can solve $\mathsf{NC}^{1}$-complete problems (McKenzie et al., 1991) including the $A_{5}$ word problem.
  2. 2.
@@
    S4, Mamba, and transformers require greater depth even $A_{4}\times\mathbb{Z}_{5}$, which can be theoretically expressed by $\mathsf{TC}^{0}$ circuits. Although transformer and Mamba models of a given depth perform as good or better on $A_{4}\times\mathbb{Z}_{5}$ as they on $A_{5}$, they still require increasingly many layers to handle proportionally longer sequences. There are two possible interpretations of this. First, it could be that while these word problems are expressible in $\mathsf{TC}^{0}$, they cannot be expressed by S4, Mamba, or transformers (which can each likely recognize only a proper subset of $\mathsf{TC}^{0}$). On the other hand, it is possible that these word problems are expressible by transformers, S4, and Mamba but that effectively learning a constant-depth solution is difficult.
  3. 3.
@@
    Despite this limitation, S4 and Mamba appear empirically better at approximate state tracking on the non-commutative tasks than the transformer. For length-$n$ sequences from $A_{4}\times\mathbb{Z}_{5}$ or $A_{5}$, the transformer requires at least as many (and frequently fewer) layers as S4 or Mamba to solve the task.
@@ 7 Conclusion
We have shown that SSMs, like transformers, can only express computation in the complexity class $\mathsf{L}$-uniform $\mathsf{TC}^{0}$. This means they cannot solve inherently sequential problems like graph connectivity, boolean formula evaluation, and—of particular interest for state tracking—the permutation composition problem $S_{5}$. $S_{5}$ can be naturally expressed by true recurrent models like RNNs and captures the essence of hard state tracking due to its $\mathsf{NC}^{1}$-completeness. In practice, one-layer RNNs can easily learn a task capturing $S_{5}$ SSMs require depth growing with the sequence length. These results reveal that SSMs cannot truly track state: rather, they can only solve simple state-tracking problems for which shallow shortcuts exist (Liu et al., 2023).
We also showed that simple extensions of SSMs can express $S_{5}$, although this comes with potential drawbacks as far as parallelism and learning dynamics. In future work, it would be interesting to more thoroughly explore the practical viability of our SSM extensions. Ultimately, this line of work has the potential to unlock new neural architectures that balance the parallelism of transformers and SSMs with full expressive power for state tracking, enabling LLMs that can benefit from scale while enjoying a greater capacity to reason about games, code, and language.
Broader Impact
@@ Appendix A Floating-Point Arithmetic
Definition A.1 (Iterated $\mathbb{D}$-product).
For $\phi_{1},\ldots,\phi_{z}\in\mathbb{D}$, their iterated $\mathbb{D}$-product is:
$\bigotimes_{i=1}^{z}\phi_{i}\,\triangleq\,\mathsf{cast}_{\mathbb{D}}\left(\prod_{i=1}^{z}\mathsf{cast}_{\mathbb{R}}(\phi_{i})\right).$ (8)
Here $\mathsf{cast}_{\mathbb{R}}$ converts a number in $\mathbb{D}$ to the corresponding real number. $\mathbb{D}$ is implicit in the notations $\mathsf{cast}_{\mathbb{R}}$ and $\bigotimes$.

### v2

@@ 4.4 Discussion
Theorems 4.2 and 4.4 establish that common SSM variants, like transformers, can only express solutions to problems in the class $\mathsf{TC}^{0}$. This means these SSMs cannot solve $\mathsf{NC}^{1}$-hard problems like evaluating boolean formulas or graph connectivity. In particular, it shows that they are limited as far as their state tracking capabilities as they are unable to compose permutations (solve the $S_{5}$ word problem):
@@ 5 Extending the Expressive Power of SSMs
We have shown that S4 and S6, despite their seemingly “stateful” design, cannot express problems outside $\mathsf{TC}^{0}$, which includes state tracking like $S_{5}$. We show how SSMs can be extended to close the gap in expressive power with RNNs, allowing them to express $S_{5}$. Two simple extensions can bring about this increase in expressive power, assuming layer input dimension $k>1$. First, adding a nonlinearity makes the SSM into an RNN, adding expressive power but degrading parallelism. On the other hand, allowing $\bar{\mathbf{A}}_{i}$ to be input-dependent makes the SSM more like a weighted finite automaton (WFA; Mohri, 2009), adding expressive power while remaining parallelizable.
@@ 5.1 Via Nonlinearities
One extension to the SSM is to add a nonlinearity, effectively making it an RNN. We call this an RNN-SSM layer:
$\mathbf{h}_{i}=\mathrm{sgn}\left(\bar{\mathbf{A}}\mathbf{h}_{i-1}+\bar{\mathbf{B}}x_{i}\right).$
A model with this architecture can solve the $S_{5}$ word problem when the input dimension $k>1$:
@@
For any regular language $L\subseteq\Sigma^{*}$ (including the word problem for $S_{5}$), there exists a one-layer log-precision RNN-SSM with $k=\left\lvert\Sigma\right\rvert$ that recognizes $L$.
@@
The standard constructions for simulating automata with RNNs (cf. Minsky, 1954; Merrill, 2019) apply. The condition $k=\left\lvert\Sigma\right\rvert$ comes from needing to represent token types with linearly independent vectors. ∎
Adding a nonlinearity to the output of an SSM layer (as in Mamba) is not the same thing as an RNN-SSM. Rather, an RNN-SSM applies the nonlinearity at each recurrent update. A downside of this approach is that it becomes nonlinear to parallelize the RNN-SSM computation graph with the SCAN algorithm used by linear SSMs (Blelloch, 1990).
@@ 5.2 Via Input-Dependent Transition Matrices
Another way to get greater expressive power is to let the transition matrix $\bar{\mathbf{A}}_{i}$ be fully input-dependent, as explored by Liquid S4 (Hasani et al., 2023). To illustrate this, we define a minimally different SSM called Input-Dependent S4 (IDS4) that achieves greater expressive power for state tracking. Let $\pi_{\mathbf{A}}:\mathbb{R}^{k}\to\mathbb{R}^{d\times d}$ be some affine transformation where the output vector is interpreted as a $d\times d$ matrix, and let $\bar{\mathbf{A}}_{i}=\pi_{\mathbf{A}}(\mathbf{x}_{i})$. Let $\bar{\mathbf{B}},\mathbf{C},\mathbf{D}$ be fixed (w.r.t. $i$). By Definition 2.1, the IDS4 convolutional form computes an iterated product of non-diagonal, input-dependent matrices:
$\mathbf{h}_{i}=\sum_{j=1}^{i}\left(\prod_{k=j+1}^{i}\pi_{\mathbf{A}}(\mathbf{x}_{i})\right)\bar{\mathbf{B}}\mathbf{x}_{j}.$
In contrast to matrix powers or iterated products of diagonal matrices, iterated products of general matrices cannot be computed in $\mathsf{TC}^{0}$ (Mereghetti & Palano, 2000). This means that the arguments from Theorems 4.2 and 4.4 will not go through for IDS4. In fact, we can show IDS4 gains expressive power beyond $\mathsf{TC}^{0}$:
Figure 3: Minimum number of layers (lower is better) required to attain $>90\%$ validation accuracy on group multiplication problems by sequence length and group. RNN and IDS4 models of constant depth can solve arbitrarily long sequences, while transformer, S4, and Mamba models require depths monotonically increasing in sequence length.
@@
For any regular language $L\subseteq\Sigma^{*}$ (including the word problem for $S_{5}$), there exists a one-layer log-precision IDS4 SSM with $k=\left\lvert\Sigma\right\rvert$ that recognizes \$$\$L$, where \$$\$\not\in\Sigma$ is a special beginning-of-string symbol.
@@
It suffices to show that IDS4 can simulate a deterministic finite automaton (DFA). We do this via a transition monoid construction. For any $w\in\Sigma^{*}$, let $\delta_{w}:Q\to Q$ be the function mapping a state to its eventual destination state after $w$ is read from that state. For any DFA, this set of functions forms a finite monoid (the transition monoid) under composition, following from the Myhill-Nerode theorem (Hopcroft et al., 2001). Further, each monoid element $\delta_{w}$ can be represented as a boolean transition matrix, making matrix multiplication isomorphic to monoid composition. Computing the transition monoid of a DFA allows recognizing valid words: compute the monoid element for a word by multiplying the elements for its tokens and then check whether the initial state maps to an accepting state.
Fix a DFA and its transition monoid $\delta$. To complete the proof, we show there exists an SSM that, for all $w\in\Sigma^{*}$, computes $\delta_{w}$ given input $x=\$w$. Let $\bar{\mathbf{A}}_{i}$ be the transition matrix representation of $\delta_{x_{i}}$. Matrix multiplication is isomorphic to composition of transition monoid elements. We view indices in $\mathbf{h}_{i}$ as states and define $\bar{\mathbf{B}}\$$ as $1$ at the initial state $q_{0}$ and 00 elsewhere. For other $\sigma$, let $\bar{\mathbf{B}}\sigma=\vec{0}$. This yields the following convolutional form:
[…]
Since $x=\$w$, we conclude that $\mathbf{h}_{\left\lvert x\right\rvert}\equiv\delta_{w}(q_{0})$. ∎
@@ 5.3 Discussion
Parallelism. To be used effectively in an LLM, a model architecture must be parallelizable on practical hardware. Architectures in $\mathsf{TC}^{0}$ are parallelizable by design (Merrill & Sabharwal, 2023a), but architectures in $\mathsf{NC}^{1}$ may still be parallelizable to log depth even if they cannot be parallelized to constant depth. For IDS4, the bottleneck would be computing iterated matrix product with a log-depth computation graph. This could be achieved with the SCAN algorithm (Blelloch, 1990) similar to S4 and S6. In contrast, it is less clear how to parallelize a model with a nonlinearity.
Learning Dynamics. Another potential concern for IDS4 is that learning dynamics could be degraded. In particular, an iterated product of matrices may lead to vanishing or exploding gradients. However, this is already potentially an issue for the S6 architecture, where the selective gating involves computing an iterated product of scalars.
@@ 6 Can SSMs Learn Permutations in Practice?
Having established theoretical limitations of SSMs for state tracking, we empirically test how well SSMs can learn such tasks, focusing on the $A_{5}$ word problem. Since this problem is $\mathsf{NC}^{1}$-complete and transformers, S4, and Mamba can only express functions in $\mathsf{TC}^{0}$, these models should require a depth that grows with the input length to solve this problem.
@@
Task. We model word problems (see Section 3.1) as a token-tagging task. Models are given as input a sequence $g_{0}g_{1}g_{2}\cdots g_{n}$ drawn from one of $A_{5}$, $A_{4}\times\mathbb{Z}_{5}$, or $\mathbb{Z}_{60}$. At each step $i$, the label is the product of the first $i$ elements of the sequence. Modeling the problem as a tagging task rather than as a sequence classification task provides the models with more supervision during training, making it as easy as possible to learn the correct function. We tokenize inputs such that each element gets a unique token.
Models. We train a transformer as a $\mathsf{TC}^{0}$ baseline, an RNN that we expect can perform state tracking, and three SSMs: S4 (Gu et al., 2022a), Mamba (Gu & Dao, 2023), and IDS4 (Section 5.2). For IDS4, we initialize the affine projection $\alpha$ as a random normal centered around the identity: $\alpha(\mathbf{x}_{i})\sim\mathbf{I}+\mathcal{N}(0,\sigma^{2})$. This ensures that, at initialization, input-dependent transitions tend to propagate the previous state, which we expect to aid learning efficiency.
Experimental Setup. We train models on sequences of length $n$ for successively larger values of $n$ and report full-sequence accuracy on a test set. [^9] We always include all $3600$ pairwise sequences of length $2$ in the training data along with the training split of length-$n$ sequences. To validate the prediction that SSMs and transformers require growing depth to solve longer $A_{5}$ word problems, we plot the minimum depth with 90% test accuracy as a function of input sequence length.
@@
Results. Figure 3 shows single-layer RNN and IDS4 models learn the word problem for arbitrarily long sequences for all three groups. In contrast, transformer, S4, and Mamba models require depth monotonically increasing in sequence length to attain good test accuracy for the non-commutative groups. We draw three conclusions from this:
@@
1. As expected, S4 and Mamba show the same limitations as transformers on the $A_{5}$ word problem. Longer $A_{5}$ sequences require deeper models, consistent with these models being in $\mathsf{TC}^{0}$. In contrast, RNNs (Theorem 5.1) and IDS4 (Theorem 5.2) can efficiently solve the $A_{5}$ word problem.
@@
2. Transformers, S4, and Mamba require greater depth even for $A_{4}\times\mathbb{Z}_{5}$, which can be theoretically expressed by $\mathsf{TC}^{0}$ circuits. Although transformer and Mamba models of a given depth perform as good or better on $A_{4}\times\mathbb{Z}_{5}$ as they on $A_{5}$, they still require increasingly many layers to handle proportionally longer sequences. There are two possible interpretations of this. First, it could be that while these word problems are expressible in $\mathsf{TC}^{0}$, they cannot be expressed by S4, Mamba, or transformers (which can each likely recognize only a proper subset of $\mathsf{TC}^{0}$). On the other hand, it is possible that these word problems are expressible by transformers, S4, and Mamba but that effectively learning a constant-depth solution is difficult.
@@
3. Despite this limitation, S4 and Mamba appear empirically better than transformer at approximate state tracking on the non-commutative tasks. For length-$n$ sequences from $A_{4}\times\mathbb{Z}_{5}$ or $A_{5}$, the transformer requires at least as many (and frequently more) layers as S4 or Mamba.
@@ 7 Conclusion
We formally analyzed a family of generalized linear SSMs and showed that, like transformers, common SSM variants including S4 and Mamba can only express computation within the complexity class $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ of highly parallel computations. This means they cannot solve inherently sequential problems like graph connectivity, boolean formula evaluation, and—of particular interest for state tracking—the permutation composition problem $S_{5}$. $S_{5}$ can be naturally expressed by true recurrent models like RNNs and captures the essence of hard state tracking due to its $\mathsf{NC}^{1}$-completeness. In practice, one-layer RNNs can easily learn a task capturing $S_{5}$ while linear SSMs require depth growing with the sequence length. These results reveal that S4, Mamba, and related SSMs cannot truly track state: rather, they can only solve simple state-tracking problems for which shallow shortcuts exist (Liu et al., 2023).
[…]
Impact Statement
@@
Acknowledgments
This work benefited from discussions with and valuable feedback from Chris Barker, Stefano Ermon, and Charles Foster. It was supported in part through the NYU IT High Performance Computing resources, services, and staff expertise. It was funded by NSF award 1922658, and WM was supported by an NSF graduate research fellowship, AI2, and Two Sigma.
@@ Appendix A Floating-Point Arithmetic
Our results use the log-precision floating point model used by Merrill & Sabharwal (2023b) to analyze transformers. For some fixed constant $c\in\mathbb{Z}^{+}$, a $c\log n$ precision float is a tuple $\langle m,e\rangle$ where $m,e$ are signed integers together taking $c\log n$ bits. Using $|x|$ to mean the number of bits used to represent integer $x$, this float represents the value $m\cdot 2^{e-|m|+1}$.
Unlike for integers, arithmetic operations over log-precision floats are not closed. That is, the product $\phi_{1}\times\phi_{2}$ of two $p$-precision floats is a well-defined number but may not be exactly representable as a $p$-precision float. It is thus necessary to define approximate versions of these operations when formalizing log-precision floating-point arithmetic. To this end, Merrill & Sabharwal (2023a) define a natural notion of approximate iterated addition over log-precision floats and show that it is computable in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$. We can naturally apply their definition of iterated addition for floats to matrices of floating points, defining iterated summation over matrices of datatype $\mathbb{D}$ as the result of treating the numbers as reals, performing exact arithmetic, and casting the exact output $\phi$ back to $\mathbb{D}$, denoted $\mathsf{cast}_{\mathbb{D}}(\phi)$. Formally:
Definition A.1 (Iterated $\mathbb{D}$-matrix sum; Merrill & Sabharwal, 2023a).
For matrices $\mathbf{M}_{1},\ldots,\mathbf{M}_{n}$ over $\mathbb{D}$ with the same size, their iterated $\mathbb{D}$-sum is
$\bigoplus_{i=1}^{z}\mathbf{M}_{i}\,\triangleq\,\mathsf{cast}_{\mathbb{D}}\left(\sum_{i=1}^{z}\mathsf{cast}_{\mathbb{R}}(\mathbf{M}_{i})\right).$
Here $\mathsf{cast}_{\mathbb{R}}$ converts a number in $\mathbb{D}$ to the corresponding real number. $\mathbb{D}$ is implicit in the notations $\mathsf{cast}_{\mathbb{R}}$ and $\bigoplus$. Integer addition can be obtained as a special case for $1$-dimensional matrices. We can also analogously defined iterated summation, which will be necessary for formalizing SSMs:
Definition A.2 (Iterated $\mathbb{D}$-matrix product).
For square matrices $\mathbf{M}_{1},\ldots,\mathbf{M}_{z}$ over $\mathbb{D}$, their iterated $\mathbb{D}$-product is
$\bigotimes_{i=1}^{z}\mathbf{M}_{i}\,\triangleq\,\mathsf{cast}_{\mathbb{D}}\left(\prod_{i=1}^{z}\mathsf{cast}_{\mathbb{R}}(\mathbf{M}_{i})\right).$
Merrill & Sabharwal (2023a) showed that iterated addition from for log-precision floats is in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$. It naturally follows from their argument that iterated addition over log-precision float matrices is also in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$. In general, iterated matrix products are not necessarily computable in $\mathsf{TC}^{0}$. However, we extend the arguments of Hesse (2001) and Mereghetti & Palano (2000) for integers to show that two special cases (iterated scalar multiplication and matrix powering) over log-precision floats are also computable in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$.
Finally, we define a canonical value for a compositional arithmetic expression over floats that enjoys the associative property.
Definition A.3 (Flattened expression evaluation).
Let $\phi$ be a compositional expression over floats, which may contain alternating sums and products as well as other operations like $\exp$. We define the canonical value of $\phi$ as the value returned by the computation graph obtained by flattening all adjacent sums into a single sum (and analogously for products).
Definition A.3 has the nice effect of making Definition A.2 associative. The only results that rely on this assumption are our analysis of diagonalizable SSMs in Appendix C. We also deal with the details of this assumption in Lemma 4.1, though the proof there also goes through directly without handling these details.
A.1 Complexity of Iterated Scalar Multiplication
The first special case of iterated matrix products we analyze is when the matrices are simply scalars (or, w.l.o.g., diagonal matrices). In this case, the iterated product can be computed in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$.